# Making it to First: The Random Access Problem in DNA Storage

### v1

@@ Abstract
We study the Random Access Problem in DNA storage, which addresses the challenge of retrieving a specific information strand from a DNA-based storage system. Given that $k$ information strands, representing the data, are encoded into $n$ strands using a code. The goal under this paradigm is to identify and analyze codes that minimize the expected number of reads required to retrieve any of the $k$ information strand, while in each read one of the $n$ encoded strands is read uniformly at random. We fully solve the case when $k=2$, showing that the best possible code attains a random access expectation of $0.914\cdot 2$. Moreover, we generalize a construction from [17], specific to $k=3$, for any value of $k$. Our construction uses $B_{k-1}$ sequences over $\mathbb{Z}_{q-1}$, that always exist over large finite fields. For $k=4$, we show that this generalized construction outperforms all previous constructions in terms of reducing the random access expectation.
@@ 1 Introduction
While several studies have demonstrated the significant potential of DNA as a data storage medium [2, 6, 7, 22, 27, 28, 3], its adoption as a practical alternative to current storage technologies is still limited by challenges related to cost and efficiency. A key factor contributing to these challenges is the coverage depth of DNA storage, defined as the ratio between the number of sequenced reads and the number of designed strands[18]. Reducing the coverage depth is critical for improving the latency of existing DNA storage systems and significantly lowering their costs. One very natural question concerning the coverage depth has been studied recently in [4] and deals with the problem of reducing the expected number of samples needed to retrieve all the stored information. This problem is closely related to the Coupon Collector’s Problem [13, 14, 12, 21], for more details, see the last paragraph in Section 2.
@@
In this work, we study the Random Access Problem, first introduced in [4], which addresses the challenge of retrieving a single specific information strand from a DNA-based storage system. Given that $k$ information strands, representing the data, are encoded into $n$ strands using an error-correcting code, the goal is to identify and analyze codes that minimize the expected number of reads required to retrieve any of the $k$ information strands, while in each read one of the $n$ encoded strands is read uniformly at random. Initial steps towards solving this problem were made in [4] and [16], where constructions, analysis of well-known codes, and bounds on the random access expectation for arbitrary codes were presented. In a later work [17], this setting was explored from a geometric perspective, leading to a construction that outperformed all previously known codes for $k=3$. Despite the progress made in these works, many essential questions remain unsolved.
@@
In this paper, we address some of the open questions in the random access setting. In particular, we provide a complete solution for the $k=2$ case, showing what the smallest possible random access expectation is in this case. Moreover, we generalize the construction from [17] for arbitrary values of $k$ and show the resulting construction outperforms all current known codes for $k=4$. The rest of the paper is organized as follows. Section 2 introduces the relevant definitions and formalizes the problem statement. Section 3 provides a solution to the random access problem for $k=2$. In Section 4, we provide a construction for general $k$ and then analyze it for $k=3$ and $k=4$. Finally, Section 5 concludes the paper and outlines several directions for future research.
@@ 2 Preliminaries
Throughout this paper, we adopt the following notations. Let $k$ and $n$ be positive integers with $k\leq n$, and let $q$ be a prime power. Denote by $[k]$ the set $\{1,2,\dots,k\}$ and by $H_{n}$ the $n$-th harmonic number, i.e., $H_{n}:=1+1/2+\dots+1/n$. Let $\mathbb{F}_{q}$ denote the finite field with $q$ elements, and let ${\mathbb{F}_{q}^{k}}$ denote the $k$-dimensional vector space over ${\mathbb{F}}_{q}$. Let ${{\boldsymbol{e}}}_{i}\in{\mathbb{F}}_{q}^{k}$ denote the $i$-th standard vector, and for a set of vectors ${\bf g}_{1},{\bf g}_{2},\dots,{\bf g}_{i}\in{\mathbb{F}}_{q}^{k}$, let $\langle{\bf g}_{1},{\bf g}_{2},\dots,{\bf g}_{i}\rangle$ denote their span.
@@
The encoding process utilizes a $k$-dimensional linear block code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$, mapping an information vector ${\boldsymbol{x}}=(x_{1},\dots,x_{k})\in\mathbb{F}_{q}^{k}$ to an encoded vector ${\boldsymbol{y}}=(y_{1},\dots,y_{n})\in\mathbb{F}_{q}^{n}$. To retrieve the stored information, the encoded strands are first synthesized and then sequenced using DNA sequencing technology. This process produces multiple erroneous copies of the strands, referred to as reads. In line with prior works [4, 16, 17], we assume that no errors are introduced during synthesis or sequencing. The output of the sequencing process is thus a multiset of unordered reads. Given the high cost and relatively low throughput of current DNA sequencing technologies compared to other archival storage systems, reducing the coverage depth—the ratio of sequenced reads to the number of encoded strands—is crucial for improving efficiency.
In this paper, we focus on the random access setting, where the goal is to retrieve a single information strand $x_{i}$ for $i\in[k]$. Previous works (e.g., [4, 16, 17]), have shown that appropriate coding schemes can reduce the expected sample size for recovering an information strand to below $k$. We illustrate this in Example 1, as presented in [16]. Note that we implicitly assume that we know the index of each of the strands that we are sampling. If we did not know the locations of the recovered symbols in the underlying code, we do not necessarily know how to recover information strands from sampled strands.
@@
Assume we want to store an information vector of size two, $(x_{1},x_{2})\in{\mathbb{F}}_{q}^{2}$. Without coding, the expected number of samples required to recover a specific information strand is 2, assuming that the samples are read uniformly at random. Now, suppose we encode the data using the following generator matrix $G$
$G=\begin{pmatrix}1&0&1&0&1\\ 0&1&0&1&1\\ \end{pmatrix}.$ (1)
The encoded data is stored as $(x_{1},x_{2})G=(x_{1},x_{2},x_{1},x_{2},x_{1}+x_{2})\in{\mathbb{F}}_{q}^{5}$. Using the known expectation of a geometric distribution and conditioning on the first sample, it can be shown that the expected number of samples required to recover a specific information strand for this case is approximately $1.917<2$. Here, “recovering” means reconstructing the original information strand as a linear combination of the sampled symbols. For instance, if the last two encoded symbols are sampled, the strand $x_{1}$ can be recovered as $x_{1}=x_{2}+(x_{1}+x_{2})$.
@@
To formalize the problem, we define the main problem addressed in this paper, building on the framework presented in [16].
Problem 1.
Let $G\in{\mathbb{F}}_{q}^{k\times n}$ be a rank-$k$ matrix. Suppose that the columns of $G$ are drawn uniformly at random, meaning that each column has a probability $1/n$ of being drawn and columns can be drawn multiple times. For $i\in[k]$, let $\tau_{i}(G)$ denote the random variable that counts the minimum number of columns of $G$ that are drawn until the standard basis vector ${{\boldsymbol{e}}}_{i}\in{\mathbb{F}}_{q}^{k}$ is in their ${\mathbb{F}}_{q}$-span. Our goals are:
@@
    Compute the expectation ${\mathbb{E}}[\tau_{i}(G)]$ and the maximum expectation:
    $T_{\max}(G)=\max_{i\in[k]}{\mathbb{E}}[\tau_{i}(G)].$
@@
    $T(k)=\liminf_{q\longrightarrow\infty}T_{q}(k).$
Note that the values of $T_{q}(n,k)$ and $T_{q}(k)$ are defined only when $q$ is a prime power. If $Q$ is the set of all prime powers, we omit writing $q\in Q$ when writing $q\to+\infty$.
The study of the values of $T_{\max}(G),\text{\leavevmode\nobreak\ }T_{q}(n,k),\text{\leavevmode\nobreak\ }T_{q}(k),$ and $T(k)$ was initiated in [4]. It was shown that for several families of codes, such as the identity code, the simple parity code, and MDS codes, it holds that ${\mathbb{E}}[\tau_{i}(G)]=k$ for every $i\in[k]$, when $G$ is a systematic generator matrix of these codes. In particular, the result for identity codes established that $T_{q}(k,k)=k$. However, determining $T_{q}(n,k)$ for general parameters remains an open and intriguing question. Initial progress towards addressing this problem was made in [4], where several constructions were proposed that achieve a maximum expectation strictly less than $k$. In particular, it was shown that $T(k=2)\leq 0.914\cdot 2,\text{\leavevmode\nobreak\ }T(k=3)\leq 0.89\cdot 3$ and for arbitrary $k$ which is a multiple of $4$, it holds that for large enough $q$: $T_{q}(n=2k,k)\leq 0.95k$. In addition, two lower bounds on $T_{q}(n,k)$ were derived in [4]: for any $n,k$ and $q$, we have that $T_{q}(n,k)\geq n-\frac{n(n-k)}{n}(H_{n}-H_{n-k})$ and $T_{q}(n,k)\geq\frac{k+1}{2}$.
@@
Using this result, it was derived that various families of codes, e.g. Hamming, simplex, systematic MDS, all achieve a random access expectation of $k$. Then in [17], by looking at the random access problem from a geometric point of view, the authors proposed a construction for the case of $k=3$. Using the above formula, they demonstrated that their construction achieves an expectation upper bounded by $0.88\overline{22}\cdot 3$, improving the result of [4]. In addition, [17] resolved a conjecture proposed in [4] regarding a class of codes with rate $1/2$, which achieves a random access expectation strictly smaller than $0.9456k$ as $k$ tends to infinity.
Despite these valuable contributions, the fundamental limits of the random access coverage depth problem are still not well understood. Even for the case of $k=2$, the values of $T_{q}(k)$ and $T(k)$ had not been determined previously. In this work, we seek to deepen our understanding of the random access problem and provide new insights into the values of $T_{\max}(G),\text{\leavevmode\nobreak\ }T_{q}(n,k),\text{\leavevmode\nobreak\ }T_{q}(k)$ and $T(k)$.
We would like to note that a closely related problem is the non-random access setting, where the goal is to retrieve all the user’s information. By drawing a connection to the Coupon Collector’s Problem [13, 14, 12, 21], the authors in [4] established that if the $k$ information strands are encoded by an MDS code, then the expected number of reads to decode all $k$ information strands is $H_{n}-H_{n-k}$. Moreover, the authors showed that one cannot do better than this, and so MDS codes are the most suitable codes for minimizing the expected number of reads when the user wants to retrieve all their data. Later on, the non-random access version of the DNA coverage depth problem was expanded in [9] to address composite DNA letters [2], and further explored in [23, 26] for the setup of the combinatorial composite of DNA shortmers [24]. A similar concept was examined in [8], where the authors analyzed the trade-offs between reading costs, tied directly to coverage depth, and writing costs. Another extension to the random access setup was studied in [1]. However, their study focused on decoding groups of DNA strands collectively representing a single file, rather than decoding individual strands.

### v2

@@ Abstract
In this paper, we study the Random Access Problem in DNA storage, which addresses the challenge of retrieving a specific information strand from a DNA-based storage system. In this framework, the data is represented by $k$ information strands which represent the data and are encoded into $n$ strands using a linear code. Then, each sequencing read returns one encoded strand which is chosen uniformly at random. The goal under this paradigm is to design codes that minimize the expected number of reads required to recover an arbitrary information strand. We fully solve the case when $k=2$, showing that the best possible code attains a random access expectation of $1+\frac{2}{\sqrt{2}+1}\approx 0.914\cdot 2$ for $q$ large enough. Moreover, we generalize a construction from [19], specifically to $k=3$, for any value of $k$. Our construction uses $B_{k-1}$ sequences over $\mathbb{Z}_{q-1}$, that always exist over large finite fields. We show that for every $k\geq 4$, this generalized construction outperforms all previous constructions in terms of reducing the random access expectation.
@@ 1 Introduction
While several studies have demonstrated the significant potential of DNA as a data storage medium [2, 8, 9, 24, 29, 30, 3], its adoption as a practical alternative to current storage technologies is still limited by challenges related to cost and efficiency. A key factor contributing to these challenges is the coverage depth of DNA storage, defined as the ratio between the number of sequenced reads and the number of designed strands[20]. Reducing the coverage depth is critical for improving the latency of existing DNA storage systems and significantly lowering their costs.
In a recent paper [4], Bar-Lev et al. initiated the study of the DNA Coverage Depth Problem, which aims to reduce the cost and latency of DNA sequencing by analyzing the expected number of reads required to retrieve the user’s information. In this model, $k$ information strands, representing the user’s data, are encoded into $n$ strands using a matrix $G\in\mathbb{F}_{q}^{k\times n}$, often refereed to as a generator matrix of an error-correcting code. The columns of $G$ are then sampled uniformly at random with repetition until the desired information is successfully retrieved (every such sample is referred to as read). The authors considered two variants of this problem: the non-random access setting and the random access setting. In the non-random access setting, the objective is to recover all $k$ information strands, i.e., the entire user’s information. In contrast, the random access setting focuses on retrieving a single, specific strand out of the $k$ information strands. We also remark that a subsequent work [1] studied a related model in which groups of strands that collectively represent a single file are decoded, rather than individual strands or the entire dataset.
By drawing a connection to the classical Coupon Collector’s Problem [15, 16, 14, 23], the authors of [4] showed that when the $k$ information strands are encoded into $n$ strands using an MDS code, then the expected number of reads required to recover all $k$ information strands is $H_{n}-H_{n-k}$, where $H_{i}$ is the $i$-th harmonic number. They further proved that no coding scheme can achieve a lower expectation-making MDS codes optimal for minimizing the expected number of reads when the goal is to retrieve the entire data. The non-random access setting was later extended in [11] to accommodate composite DNA letters [2], and further explored in [25, 28] in the context of combinatorial composites of DNA shortmers [26]. A related direction was examined in [10], where the authors analyzed the trade-offs between reading costs, tied directly to coverage depth, and writing costs. Finally, [5] studied the non–random-access setting for codes over small finite fields.
@@
In this work, we focus on the random access setting, which remains relatively less understood and presents many open questions. Notably, while MDS codes are optimal in the non-random access setting, they yield an expected number of reads equal to $k$ in the random access case [4] (that is, the expected number of reads required to retrieve a specific information strand is $k$). This is as good (or bad) as not using any code at all and using the identity code, which also gives a random access expectation equal to $k$. However, as demonstrated in [4], codes with strictly lower expectation do exist. Our goal is to identify and analyze such codes, i.e., those that minimize the expected number of reads required to retrieve any one of the $k$ original strands. Initial steps towards this problem were made in [4] and [18], where constructions, analysis of well-known codes, and bounds on the random access expectation for arbitrary codes were presented. In a later work [19], this setting was explored from a geometric perspective, leading to a construction that outperformed all previously known codes for $k=3$. More recently, the authors of [6] resolved a conjecture proposed in [4] concerning codes of rate $\frac{1}{2}$, and provided a closed-form expression for the random access expectation of the geometric construction for $k=3$ from [19]. Despite these advancements, many essential questions remain unsolved.
@@
In this paper, we address some of the open questions in the random access setting. In particular, we provide a complete solution for the $k=2$ case, finding the smallest possible random access expectation in this case. Moreover, we generalize the construction of [19] to arbitrary values of $k$, and show that the resulting code outperforms all current known constructions for $k\geq 4$.
The rest of the paper is organized as follows. Section 2 introduces the relevant definitions and formalizes the problem statement. Section 3 provides a solution to the random access problem for $k=2$. In Section 4, we provide a construction for general $k$ and then analyze if for all values, showing it outperforms all current constructions. Finally, Section 5 concludes the paper and outlines several directions for future research.
@@ 2 Preliminaries
Throughout this paper, we adopt the following notations. Let $k$ and $n$ be positive integers with $k\leq n$, and let $q$ be a prime power. Denote by $[k]$ the set $\{1,2,\dots,k\}$ and by $H_{n}$ the $n$-th harmonic number, i.e., $H_{n}:=1+1/2+\dots+1/n$. Let $\mathbb{F}_{q}$ denote the finite field with $q$ elements, and let $\mathbb{F}_{q}^{k}$} denote the $k$-dimensional vector space over ${\mathbb{F}}_{q}$. For $i\in[k]$ denote by ${{\boldsymbol{e}}}_{i}\in{\mathbb{F}}_{q}^{k}$ the $i$-th standard vector, and for an integer $\ell\geq 0$ and a set of vectors ${\bf g}_{1},{\bf g}_{2},\dots,{\bf g}_{\ell}\in{\mathbb{F}}_{q}^{k}$, let $\langle{\bf g}_{1},{\bf g}_{2},\dots,{\bf g}_{\ell}\rangle$ denote their span. For a set $S$, we denote by $\mathcal{P}r(S)$ the set of all probability distributions over $S$.
@@
The encoding process utilizes a $k$-dimensional linear block code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$, mapping an information vector ${\boldsymbol{x}}=(x_{1},\dots,x_{k})\in\mathbb{F}_{q}^{k}$ to an encoded vector ${\boldsymbol{y}}=(y_{1},\dots,y_{n})\in\mathbb{F}_{q}^{n}$. In this storage system, the encoded strands are first synthesized and then sequenced using DNA sequencing technology. This process produces multiple erroneous copies of the strands, referred to as reads. In line with prior works [4, 18, 19], we assume that no errors are introduced during synthesis or sequencing. The output of the sequencing process is thus a multiset of unordered reads of encoded strands. Given the high cost and relatively low throughput of current DNA sequencing technologies compared to other archival storage systems, reducing the coverage depth—the ratio of sequenced reads to the number of encoded strands—is crucial for improving efficiency.
In this paper, we focus on the random access setting, where the goal is to retrieve a single information strand ${\boldsymbol{x}}_{i}$ for $i\in[k]$. Previous works (e.g., [4, 6, 18, 19]), have shown that appropriate coding schemes can reduce the expected sample size for recovering an information strand to below $k$. We illustrate this in Example 1, as presented in [18].
@@
Assume we want to store an information vector of length two, $(x_{1},x_{2})\in{\mathbb{F}}_{q}^{2}$. Without coding, the expected number of samples required to recover a specific information strand is 2, assuming that the samples are read uniformly at random. Now, suppose we encode the data using the following generator matrix:
$G=\begin{pmatrix}1&0&1&0&1\\ 0&1&0&1&1\\ \end{pmatrix}\in{\mathbb{F}}_{q}^{2\times 5}.$ (1)
The encoded data is stored as $(x_{1},x_{2})G=(x_{1},x_{2},x_{1},x_{2},x_{1}+x_{2})\in{\mathbb{F}}_{q}^{5}$. It can be shown (see Claim 1 in Section 3) that the expected number of samples required to recover a specific information strand for this case is approximately $1.917<2$. Here, “recovering” means reconstructing the original information strand as a linear combination of the sampled symbols. For instance, if the last two encoded symbols are sampled, the strand $x_{1}$ can be recovered as $x_{1}=-x_{2}+(x_{1}+x_{2})$.
@@
We now formally define the main problem addressed in this paper, building on the framework presented in [18].
Problem 1 (The Random Access Problem).
Let $G\in{\mathbb{F}}_{q}^{k\times n}$ be a rank-$k$ matrix. Suppose that the columns of $G$ are drawn uniformly at random with repetition, meaning that each column has a probability $1/n$ of being drawn and columns can be drawn multiple times. For $i\in[k]$, let $\tau_{i}(G)$ denote the random variable that counts the minimum number of columns of $G$ that are drawn, until the standard basis vector ${{\boldsymbol{e}}}_{i}\in{\mathbb{F}}_{q}^{k}$ is in their ${\mathbb{F}}_{q}$-span. Our goals are:
@@
    Construct full-rank matrices $G\in{\mathbb{F}}_{q}^{k\times n}$ for which
    $T_{\max}(G):=\max_{i\in[k]}{\mathbb{E}}[\tau_{i}(G)]<k.$
@@
    $T(k):=\liminf_{q\to\infty}T_{q}(k).$
Note that the values of $T_{q}(n,k)$ and $T_{q}(k)$ are defined only when $q$ is a prime power. If $Q$ is the set of all prime powers, we omit writing $q\in Q$ when writing $q\to+\infty$. In addition, note that if a vector $\boldsymbol{x}$} appears $n_{x}$ times as a column in $G$, for some integer $n_{x}\geq 1$, then it has a probability of $\frac{n_{x}}{n}$ to be selected at each draw. In some cases, it is more convenient to approach the problem from a probabilistic perspective. To this end, we introduce the following related problem.
Problem 2.
Let $\mathcal{P}:=\{\langle v\rangle:v\in\mathbb{F}_{q}^{\,k}\setminus\{0\}\}$ be the set of one–dimensional subspaces of $\mathbb{F}_{q}^{\,k}$, that is, the projective points of PG(k1,q)PG(k-1,q). Select one non-zero representative from each subspace and arrange these vectors as the columns of a matrix $H\in{\mathbb{F}}_{q}^{k\times|\mathcal{P}|}$, so that every projective point appears exactly once. For a probability distribution $\mu$ on $\mathcal{P}$, sample the columns of $H$ with repetition according to $\mu$. For each $i\in[k]$, let $\tau_{i}(\mu)$ denote the random variable that counts the minimum number of columns of $H$ that are drawn until the standard basis vector ${{\boldsymbol{e}}}_{i}\in{\mathbb{F}}_{q}^{k}$ lies in their ${\mathbb{F}}_{q}$-span. Our goals are:
  1. 1.
    Compute the expectation ${\mathbb{E}}[\tau_{i}(\mu)]$ and the maximum expectation:
    $T_{\max}(\mu)=\max_{i\in[k]}{\mathbb{E}}[\tau_{i}(\mu)].$
  2. 2.
    Determine the smallest possible maximum random access expectation among all possible distributions:
    $\overline{T}_{q}(k)=\min_{\mu\in\mathcal{P}r{[{\mathcal{P}}]}}T_{\max}(\mu).$
Problems $1$ and $2$ are closely connected. As observed before stating Problem 2, any matrix $G\in{\mathbb{F}}_{q}^{k\times n}$ induces a probability distribution over $\mathcal{P}r{[\mathcal{P}]}$, implying that $\overline{T}_{q}(k)\leq{T}_{q}(k)$. Conversely, for any probability distribution in $\mathcal{P}r{[\mathcal{P}]}$, there exists a rational probability distribution that is arbitrarily close to it, and for large enough $n$, it can be realized via a matrix $G\in{\mathbb{F}}_{q}^{k\times n}$, hence we have that $\overline{T}_{q}(k)={T}_{q}(k)$.
The study of the values of $T_{\max}(G),\text{\leavevmode\nobreak\ }T_{q}(n,k),\text{\leavevmode\nobreak\ }T_{q}(k),$ and $T(k)$ was initiated in [4]. It was shown that for several families of codes, such as the identity code, the simple parity code, and MDS codes, it holds that ${\mathbb{E}}[\tau_{i}(G)]=k$ for every $i\in[k]$, when $G$ is a systematic generator matrix of these codes. In particular, the result for identity codes established that $T_{q}(k,k)=k$. However, determining $T_{q}(n,k)$ for general parameters remains an open and intriguing question.
Initial progress towards addressing this problem was made in [4], which introduced several code families whose random-access expectation is strictly below $k$. In particular, it was shown that $T(k=2)\leq 0.914\cdot 2,\text{\leavevmode\nobreak\ }T(k=3)\leq 0.89\cdot 3$ and for arbitrary $k$ which is a multiple of $4$, it holds that for large enough $q$: $T_{q}(n=2k,k)\leq 0.95k$. The authors of [4] also conjectured that their rate-$1/2$ codes have a random-access expectation that drops below $0.9456k$ as $k\to\infty$. This conjecture was later resolved in [19], and the precise asymptotic limit was established in [6], proving that $\lim_{k\to\infty}\frac{T_{q}(n=2k,k)}{k}\leq\frac{8\sqrt{3}\pi-18}{27}$. Moreover, [4] proved that for all positive integers $a$ and $k$, it holds that $\frac{T(ak)}{ak}\leq\frac{T(k)}{k}$, showing that the per-strand expectation does not increase when the message length is scaled. In addition, two lower bounds on $T_{q}(n,k)$ were derived in [4]: $T_{q}(n,k)\geq n-\frac{n(n-k)}{n}(H_{n}-H_{n-k})$ and $T_{q}(n,k)\geq\frac{k+1}{2}$.
@@
Using this result, it was derived that for various families of codes, e.g. Hamming, simplex, systematic MDS, Golay, Reed-Muller, the random access expectation is $k$. Then, in [19], by looking at the random access problem from a geometric point of view, the authors proposed a construction for the case of $k=3$. Using the above formula, they demonstrated that their construction achieves a maximal random-access expectation upper bounded by $0.88\overline{22}\cdot 3$, improving the result of [4]. Later, the authors of [6] computed the exact expectation of this geometric construction and, with optimal parameters, obtained a random access expectation of about $0.881542\cdot 3$.
Despite these valuable contributions, the fundamental limits of the random access coverage depth problem are still not well understood. Even for the case of $k=2$, the values of $T_{q}(k)$ and $T(k)$ had not been determined previously. In this work, we seek to deepen and extend the knowledge of the random access problem and provide new insights into the values of $T_{\max}(G),\text{\leavevmode\nobreak\ }T_{q}(n,k),\text{\leavevmode\nobreak\ }T_{q}(k)$ and $T(k)$.